If cost of meat per kg is 240 per kg and shopkeeper provides discount of 15% on every 3kg. What will be cost of 4 kg of meat?
Found 3 solutions by Alan3354, Shin123, richwmiller:
Answer by Alan3354(69443)
Answer by Shin123(626)
Answer by richwmiller(17219)

The unclear part is how much is charged for the 4th kg.
Does the 4th kg get the 15% discount or is it charged 240 or something else?
The first three kg would cost (3*240)*.85=612
If charging 240 for the 4th kg then the total is 612+240=852
If the 4th kg gets the 15% discount then the total is (4*240)*.85=816
if something else, then who knows.

Warning: I am still sick in bed. Double check my work.